OpenStudy (anonymous):

Solve 1n 2 + 1n x = 5. Round to the nearest thousandth, if necessary.

OpenStudy (anonymous):

Use laws of logarithms to combine left side ln(2x)=5 do base e to both sides 2x=e^5 divide by 2 x=(e^5)/2

OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

You can use a calculator to get decimal answer

OpenStudy (anonymous):

ho do u divide a variable by a number?

OpenStudy (anonymous):

e is eulers number, its about 2.718 i think. There should be an e on your calculator if you have a good one

OpenStudy (anonymous):

i got 74.20657955?

OpenStudy (anonymous):

Thats right

OpenStudy (anonymous):

thx!